### v1

# Deep learning-based sequential data assimilation for chaotic dynamics identifies local instabilities from single state forecasts

Abstract
We investigate the ability to discover data assimilation (DA) schemes meant for chaotic dynamics with deep learning (DL). The focus is on learning the analysis step of sequential DA, from state trajectories and their observations, using a simple residual convolutional neural network, while assuming the dynamics to be known. Experiments are performed with the Lorenz 96 dynamics, which display spatiotemporal chaos and for which solid benchmarks for DA performance exist. The accuracy of the states obtained from the learned analysis approaches that of the best possibly tuned ensemble Kalman filter (EnKF), and is far better than that of variational DA alternatives. Critically, this can be achieved while propagating even just a single state in the forecast step. We investigate the reason for achieving ensemble filtering accuracy without an ensemble. We diagnose that the analysis scheme actually identifies key dynamical perturbations, mildly aligned with the unstable subspace, from the forecast state alone, without any ensemble-based covariances representation. This reveals that the analysis scheme has learned some multiplicative ergodic theorem associated to the DA process seen as a non-autonomous random dynamical system.
Data assimilation (DA) estimates the state of dynamical systems from sparse and noisy observations, and is used worldwide in numerical weather prediction centers. Accurate DA demands the representation of the time-dependent errors in this state estimate, usually achieved through the propagation of an ensemble of states. Using deep learning, we discover the update step of DA applied to chaotic dynamics. We show that a simple convolutional neural network (CNN) can learn DA, reaching an accuracy as good as that of ensemble-based DA. Crucially, the CNN can achieve this best accuracy with single state forecasts. This is explained by the CNN’s ability to identify local space patterns from this one state, which are used to assess the errors in the analysis. This suggests building a new class of efficient DL-based ensemble-free DA algorithms.
I Introduction
I.1 Sequential data assimilation and accuracy
Sequential data assimilation (DA) is a set of algorithms required to accurately track the trajectory of chaotic dynamics, using a dynamical model and imperfect observations. It is widely used in numerical weather prediction (NWP), and in many areas of climate sciences, (Carrassi et al., 2018) as a suite of both research and operational tools. Classical DA methods are classified into (i) variational methods, such as 3D-Var and 4D-Var, (ii) ensemble-based statistical methods, such as the ensemble Kalman filter (EnKF), and (iii) ensemble variational methods which inherit the assets of the two previous categories.(Asch, Bocquet, and Nodet, 2016) On the one hand, variational methods account for the nonlinearity of models (dynamical model and observation operators), leveraging nonlinear optimization techniques. Ensemble-based methods on the other hand can capture the errors of the day, i.e. time-dependent error statistics, via an ensemble meant to diagnose sample error statistics. Those are key properties that drive the performance of these DA methods in mildly nonlinear chaotic models. For low-order, chaotic dynamics such as the celebrated Lorenz 96 (L96) model,(Lorenz and Emanuel, 1998) the EnKF significantly outperforms 3D-Var, or a moderately-long window 4D-Var in terms of accuracy, owing to its dynamical representation of the errors. This has been emphasized and illustrated in twin experiments. (Bocquet and Sakov, 2013) In fact, current implementations of 4D-Var in NWP centers, incorporate a forecast ensemble so as to capture the errors-of-the-day.(Raynaud, Berre, and Desroziers, 2012; Bonavita, Isaksen, and Hólm, 2012) However, in high-dimensional models, these ensemble-based error statistics must necessarily be regularized using techniques known as localization and possibly inflation.(Evensen, 2009)
I.2 Time-dependent error statistics and dynamical systems
With a focus on the time-dependent error statistics of sequential DA, it has been conjectured (Carrassi et al., 2008a; Palatella, Carrassi, and Trevisan, 2013) then proven (Gurumoorthy et al., 2017; Bocquet et al., 2017; Crisan and Ghil, 2023) that for linear dynamics and when localization is unnecessary, the forecast and analysis error covariance matrices of the EnKF are confined to the unstable-neutral subspace, denoted as ${\mathcal{U}}$ from now on, of the dynamics. This subspace is spanned by the covariant Lyapunov vectors associated to non-negative Lyapunov exponents.(Legras and Vautard, 1996) It is precisely when the ensemble size is smaller than the dimension of this subspace that localization is required to avoid divergence of the EnKF. Deviating from linear dynamics turns those exact results into approximations, for which these findings were nonetheless numerically confirmed.(Bocquet and Carrassi, 2017; Chen, Carrassi, and Lucarini, 2021; Carrassi et al., 2022)
I.3 End-to-end data assimilation and deep learning
This paper focuses on methodological DA and on what can be discovered from deep learning (DL) techniques to improve state-of-the-art DA schemes such as those mentioned above. Hence, we hereby give a brief account on the recent introduction of DL techniques for DA applied to chaotic dynamics.(Cheng et al., 2023)
It was first proposed to learn DA analysis through DL from the data produced by existing DA schemes. (Härter and de Campos Velho, 2012; Cintra and de Campos Velho, 2018) One can alternatively replace the solver of a 4D-Var over a long DA window by a DL operator that would learn the outcome of the 4D-Var cost function minimization.(Fablet et al., 2021; Frerix et al., 2021; Lafon, Fablet, and Naveau, 2023; Filoche et al., 2023; Keller and Potthast, 2024) However, the latter approaches do not consider cycling sequential DA, the focus of the present paper. A systematic, formal Bayesian view on the use of DL in the critical components of sequential DA has been proposed (Boudier et al., 2023) and called data assimilation network (DAN). In the present paper, a simplified variant of this DAN concept is used. As far as ensemble and Kalman-related DA methods are concerned, it has been proposed to learn their Kalman gain, (Hoang, De Mey, and Talagrand, 1994; Hoang et al., 1998) or parameters thereof, possibly relying on an auto-differentiable implementation of the (En)KF. (Haarnoja et al., 2016; Chen, Sanz-Alonso, and Willett, 2022; Luk et al., 2024) A step further, it was also proposed to learn the fully analysis operator using (self-)supervision.(McCabe and Brown, 2021; Boudier et al., 2023) Finally, bypassing the need for dynamical models and DA schemes altogether, DL-based end-to-end methods aim at estimating states of the system from the observations only,(McNally et al., 2024; Vaughan et al., 2024) but with a focus on the feasibility of such endeavor for now.
I.4 Objective
In this paper, the forecast model is assumed to be known so as to disentangle intricate interactions of learning simultaneously the DA operators and the dynamics altogether. Our objective is to learn the analysis step of a sequential DA scheme meant for chaotic dynamics from a long trajectory of the dynamical system and the associated set of noisy, potentially partial observations. Supervised learning is used although self-supervised learning(Bocquet et al., 2019; Brajard et al., 2020; McCabe and Brown, 2021) could have been used instead but is more challenging and rather disconnected from the main results of this paper. The (incremental) analysis CNN operator will be referred to as $a_{\bm{\theta}}$, while the full resulting DA scheme will be called DAN.
We will show that the learned $a_{\bm{\theta}}$ directly discovers and utilizes a fine knowledge of the dynamics, as opposed to agnostic classical DA. The nature of these fine dynamical structures learned through $a_{\bm{\theta}}$ will then be discussed and interpreted, together with key implications for the development of modern DA methods.
II Experimental setup
We use a twin experiment setup where the states of a trajectory of the dynamics, ${\mathbf{x}}_{k}^{\mathrm{t}}\in\mathbb{R}^{N_{\mathrm{x}}}$ at equally spaced times $t_{k}$ for $k=1\ldots N_{\mathrm{c}}$, are generated from the known dynamics; the superscript “t” stands for truth. A set of observation vectors $\mathbf{y}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}}$ are obtained from these states ${\mathbf{x}}_{k}^{\mathrm{t}}$ via an observation operator $\mathcal{H}_{k}$, and perturbed by a white-in-time Gaussian noise ${\bm{\varepsilon}}_{k}$ of mean $\mathbf{0}$ and covariance matrix $\mathbf{R}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}\times N_{\mathrm{y},k}}$:
$\mathbf{y}_{k}=\mathcal{H}_{k}({\mathbf{x}}_{k}^{\mathrm{t}})+{\bm{\varepsilon}}_{k},\qquad{\bm{\varepsilon}}_{k}\sim N(\mathbf{0},\mathbf{R}_{k}).$ (1)
II.1 Analysis operator and its neural network representation
Let us define a sequential DA scheme, based on an analysis and forecast (also called background) ensemble. The $i$-th members of the analysis and forecast ensembles at time $t_{k}$ are noted ${\mathbf{x}}_{k}^{\mathrm{a},i}$ and ${\mathbf{x}}_{k}^{\mathrm{f},i}$, respectively. Denoting ${\mathcal{S}}_{\mathrm{e}}=1\ldots N_{\mathrm{e}}$, the corresponding analysis and forecast ensembles are ${\mathbf{E}}_{k}^{\mathrm{a}}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, ${\mathbf{E}}_{k}^{\mathrm{f}}=\left\{{\mathbf{x}}_{k}^{\mathrm{f},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, respectively, where ${\mathcal{E}}^{\mathrm{e}}=\mathbb{R}^{N_{\mathrm{e}}\times N_{\mathrm{x}}}$. The initial ensemble ${\mathbf{E}}_{0}^{\mathrm{f}}$ is obtained from perturbing a random state on the attractor of the dynamics.
The analysis step of the DA scheme is given by the (incremental) analysis operator $a_{\bm{\theta}}$, which depends on a set of CNN weights and biases, a vector $\bm{\theta}$,
${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+a_{\bm{\theta}}\left({\mathbf{E}}_{k}^{\mathrm{f}},\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right),$ (2a)
where ${\bm{\delta}}_{k}$, the innovation at time $t_{k}$, is defined by
${\bm{\delta}}_{k}\overset{\Delta}{=}\mathbf{y}_{k}-\mathcal{H}_{k}\left({\bar{\mathbf{x}}}_{k}^{\mathrm{f}}\right),\quad{\bar{\mathbf{x}}}_{k}^{\mathrm{f}}\overset{\Delta}{=}\frac{1}{N_{\mathrm{e}}}\sum_{i\in{\mathcal{S}}_{\mathrm{e}}}{\mathbf{x}}_{k}^{\mathrm{f},i}.$ (2b)
$\mathbf{H}_{k}$ is the tangent linear operator of $\mathcal{H}_{k}$ but any arbitrary injective operator from $\mathbb{R}^{N_{\mathrm{y},k}}$ to $\mathbb{R}^{N_{\mathrm{x}}}$ could be chosen instead. The DA forecast step propagates the analysis ensemble,
${\mathbf{E}}_{k+1}^{\mathrm{f}}={\mathcal{M}}\left({\mathbf{E}}_{k}^{\mathrm{a}}\right),$ (3)
member-wise; $\mathcal{M}$ is the resolvent over $t_{t+1}-t_{k}$ of the autonomous, i.e. time-independent, dynamics.
In this paper, we choose $a_{\bm{\theta}}$ to be a simple residual CNN, whose architecture is detailed in Appendix A. The internal convolutional layers are chosen to have $N_{\mathrm{f}}$ channels. Hence, the internal state of the CNN consists of $N_{\mathrm{f}}$ copies of the latent space which we simply chose to be isomorphic to the state space $\mathbb{R}^{N_{\mathrm{x}}}$. Furthermore, the encoder and decoder from state space to latent space and back, are chosen to coincide with the identity.
II.2 Training scheme
Towards efficiently learning an optimal $a_{\bm{\theta}}$, we consider $N_{\mathrm{r}}$ such DA runs, based on as many independent concurrent trajectories of the dynamics and as many sequences of observation vectors. The DA runs are hence specified by ${\mathbf{E}}_{k}^{\mathrm{a},r}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i,r}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}$ and ${\mathbf{E}}_{k}^{\mathrm{f},r}$ for $r=1\ldots N_{\mathrm{r}}$ and, being iterates through $N_{\mathrm{c}}$ cycles, they depend on $\bm{\theta}$, except for the set of initial conditions ${\mathbf{E}}_{0}^{\mathrm{f},r}$. In order to learn an optimal $a_{\bm{\theta}}$, a loss function is defined:
${\mathcal{L}}(\bm{\theta})=\sum_{r=1}^{N_{\mathrm{r}}}\sum_{k=1}^{N_{\mathrm{c}}}\left\|{\mathbf{x}}_{k}^{\mathrm{t},r}-{\bar{\mathbf{x}}}_{k}^{\mathrm{a},r}(\bm{\theta})\right\|^{2},\quad{\bar{\mathbf{x}}}_{k}^{\mathrm{a},r}\overset{\Delta}{=}\frac{1}{N_{\mathrm{e}}}\sum_{i\in{\mathcal{S}}_{\mathrm{e}}}{\mathbf{x}}_{k}^{\mathrm{a},i,r},$ (4)
[…]
It must be pointed out that a proper sequential DA process when seen as a dynamical system, is stable. (Carrassi et al., 2008b, c) Hence, after a rough starting phase in the training, the learned $a_{\bm{\theta}}$ should yield a numerically stable prediction-assimilation dynamical system. In particular, this is likely to avoid exploding gradients. By contrast, the task of learning a DL emulator of the dynamics over many consecutive time steps fails because of the unstable nature of the chaotic dynamics.
The $N_{\mathrm{r}}$ trajectories are dispatched into a training and validation dataset with a $90\%-10\%$ ratio. Moreover, the testing dataset is generated from an independently generated trajectory, long enough to yield converged statistics. In the test stage, the DAN scheme is used within a twin experiment using the trajectories and resulting observations of the testing dataset. Its performance is assessed using the time-averaged root mean square error (aRMSE) against the truth.
III Numerical results
The $a_{\bm{\theta}}$ operator is trained on the L96 model, and the results will be interpreted and discussed in the context of this model. The L96 model is a one-dimensional model defined over a periodic band of latitude of the Earth atmosphere. Its ordinary differential equations read
$\frac{{\mathrm{d}}x_{n}}{{\mathrm{d}}t}=(x_{n+1}-x_{n-2})x_{n-1}-x_{n}+F,$ (5)
where $x_{N_{x}}=x_{0}$, $x_{-1}=x_{N_{x}-1}$, $x_{-2}=x_{N_{x}-2}$, $F=8$, and $N_{\mathrm{x}}=40$ in the basic configuration. The Lyapunov time of this model is $0.60$; it has $13$ positive Lyapunov exponents and a zero Lyapunov exponent as a continuous-in-time autonomous system. Hence, the dimension of ${\mathcal{U}}$ is $N_{\mathrm{u}}=14$.
III.1 Hyperparameters sensitivity analysis
We first carry out a large set of trainings to assess the sensitivity of $a_{\bm{\theta}}$’s performance to its hyperparameters. We chose $N_{\mathrm{iter}}=16$, without any significant gain beyond this value while the numerical cost increases due to a deeper backpropagation. We assume the model to be fully observed with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, the identity matrix in $\mathbb{R}^{N_{\mathrm{x}}}$, and the observations to be affected by a white-in-time unbiased Gaussian noise of covariance matrix $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, for all time steps. This configuration is very often used to benchmark new DA schemes. The ensemble size $N_{\mathrm{e}}$ and the number of filter $N_{\mathrm{f}}$ were selected in a set ranging between $1$ and $40$. The number $N_{\mathrm{b}}$ of residual blocks in the CNN and number of subblocks $N_{\mathrm{sb}}$ in each residual block were both chosen in the set $\llbracket 1,6\rrbracket$. Because the L96 model has short-range correlations in space, we choose a kernel size of $f_{\mathrm{v}}=5$, even though the CNN receptive field is much larger. Overfitting is prevented by an early stopping of the minimization based on the validation score.
III.2 First results and robustness
The test DA runs with the trained $a_{\bm{\theta}}$ are actually all stable in time, yielding an aRMSE significantly below $1$, as expected if DA has any skill over the mere observations.
Unsurprisingly, we found that the larger the hyperparameters $N_{\mathrm{f}},N_{\mathrm{b}},N_{\mathrm{sb}}$, the smaller the test aRMSEs of the resulting DANs, but that $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$ offer a good compromise for accuracy versus training cost and CNN size. This will be the reference configuration.
One obvious essential drawback of learning $a_{\bm{\theta}}$ is its non-universality. Specifically, $a_{\bm{\theta}}$ depends on the observation setup used in the training dataset. This is a critical research path for end-to-end DA. Although not the aim of the present paper, we nonetheless checked the performance of the trained $a_{\bm{\theta}}$, with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\sigma_{\mathrm{y}}^{2}\mathbf{I}_{\mathrm{x}}$ with $\sigma_{\mathrm{y}}=1$, in test DA runs with similar observations but generated with $\sigma_{\mathrm{y}}$ taking value in between $0.1$ and $3.2$. Yet, in all test runs, DAN remains robust with slightly degraded aRMSEs for $\sigma_{\mathrm{y}}<1$ but aRMSEs at least as good for $\sigma_{\mathrm{y}}>1$, compared to a well-tuned EnKF. Well-tuned EnKF always refers here to an EnKF with an ensemble large enough such that localization is unnecessary and relying on the EnKF-N (Bocquet, Raanes, and Hannart, 2015; Raanes, Bocquet, and Carrassi, 2019) to optimally counteract residual sampling errors such that inflation is unnecessary. We also learned a unique $a_{\bm{\theta}}$ from observation networks whose density $N_{\mathrm{y}}/N_{\mathrm{x}}$ is randomly and uniformly chosen in the interval $[0,1]$, and $\sigma_{\mathrm{y}}=1$. This DAN was then tested on several DA runs, each one with a fixed observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$ taking value in the interval $[0.2,1]$. In this configuration, $a_{\bm{\theta}}$ performs almost as well or better than well-tuned EnKFs for $0.650.35<N_{\mathrm{y}}/N_{\mathrm{x}}<$ 0.65 and is suboptimal (compared to the EnKF) but still stable outside of this range. These results already pleasantly suggest that the findings that follow may remain valid well beyond the specifications of observation operators from which $a_{\bm{\theta}}$ was learned.
III.3 One state forecast
Using the reference configuration but with an ensemble size $N_{\mathrm{e}}$ taking values in the set $\llbracket 1,40\rrbracket$, test aRMSEs fluctuate in between $0.19$ and $0.20$. By contrast, a sizable ensemble is, as we recalled in Sec. I, one of the key reason for the success of the EnKF. For quantitative comparison, we checked that 3D-Var scores an aRMSE of $0.40$, the best linear filter, i.e. $a_{\bm{\theta}}$ learned without activation functions scores $0.38$, whereas well-tuned EnKFs with $N_{\mathrm{e}}=20$ and $N_{\mathrm{e}}=40$ score $0.191$ and $0.179$, respectively. Note that the reference $a_{\bm{\theta}}$ but with $N_{\mathrm{f}}=100$ yields an aRMSE of $0.185$, closer to the best EnKF with $N_{\mathrm{e}}=N_{\mathrm{x}}=40$, showing that further improvements are possible even though not the focus of this paper.
However, the key remark is that a single state forecast, $N_{\mathrm{e}}=1$ in $a_{\bm{\theta}}$, is as efficient as using a large ensemble. Furthermore, the need for localization and inflation is completely obviated. We have checked that this is obtained concurrently to a feature collapse in $a_{\bm{\theta}}$,(van Amersfoort et al., 2022) i.e. all channels’ last layer feature maps converge to the same state. It is likely that a better local minimum of the loss could be obtained with complex encoder and decoder (Peyron et al., 2021) and infusing diversity in the NN through Monte Carlo dropouts,(McCabe and Brown, 2021) so as to obtain an $a_{\bm{\theta}}$ leveraging the ensemble. Nonetheless, the local minimum reached in our trainings, yield an accuracy with $N_{\mathrm{e}}=1$ worthy of a well-tuned EnKF. That is why we shall concentrate in the following on interpreting this astonishing result, especially in Sec. IV, using dynamical systems theory.
Consequently, the analysis operator is hereafter learned in the reference configuration but with $N_{\mathrm{e}}=1$.
IV Interpretation and supporting results
In this section, we focus on the main result: a learned DA method with a single state forecast achieves performance on par with a well tuned EnKF. We would like to understand the reason for this performance and what $a_{\bm{\theta}}$ is learning.
IV.1 Kalman gain-based expansion
Towards this goal, we build an incremental classical Kalman update (Kalman, 1960; Ghil and Malanotte-Rizzoli, 1991) that would be a good match to $a_{\bm{\theta}}$ seen as a mathematical map, at least for small analysis increments. The first diagnostic is the mean anomaly generated by $a_{\bm{\theta}}$, i.e. how much $a_{\bm{\theta}}({\mathbf{x}},\mathbf{0})$ deviates from $\mathbf{0}$ on average. It should be small since a vanishing innovation $\bm{\delta}_{k}$ should not yield any state update. Hence, we define the time-dependent normalized scalar anomalies
$b_{k}=\frac{1}{\sqrt{N_{\mathrm{x}}}}\left\|a_{\bm{\theta}}({\mathbf{x}}_{k},\mathbf{0})\right\|,$ (6)
along with the associated mean bias $b$ and the standard deviation $s$ of $b_{k}$ in time.
Next, expanding with respect to the innovation, the following functional form for $a_{\bm{\theta}}$ is assumed:
$a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta})\approx\mathbf{K}({\mathbf{x}})\cdot\bm{\delta},$ (7)
owing to the fact that no state update is needed when the innovation vanishes, and only keeping the leading order term in $\bm{\delta}$. This is an Ansatz of $a_{\bm{\theta}}$ where $\mathbf{K}({\mathbf{x}})\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$ is meant to stand as a Kalman gain surrogate. With the propagation of a single state, classical sequential DA methods would typically resemble 3D-Var, and the gain would not depend on the forecast state (the first input variable of $a_{\bm{\theta}}$). The 3D-Var equivalent gain would be
$\mathbf{K}_{\mathrm{3D-Var}}=\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\left(\mathbf{R}+\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\right)^{-1},$ (8)
where $\mathbf{R}$ and/or $\mathbf{H}$ might depend on time if the monitoring system does so, and where $\mathbf{B}$ represents some expected climatology of the DA errors.
Interestingly, we also learned $a_{\bm{\theta}}$ replacing Eq. (2a) with the simplified ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+a_{\bm{\theta}}\left(\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right)$, whereby losing $a_{\bm{\theta}}$’s ability to extract information from ${\mathbf{E}}_{k}^{\mathrm{f}}$. This yields an aRMSE of $0.38$ in test DA runs, unsurprisingly close to the $0.40$ of our 3D-Var. Hence, learning an optimal constant-in-time $\mathbf{K}$ of an (En)KF(Luk et al., 2024), a configuration subsumed by this specific $a_{\bm{\theta}}$, is significantly suboptimal in this context.
IV.2 Estimation of the Kalman gain and error covariances
Once $a_{\bm{\theta}}$ has been obtained from training, and considering a fixed background state ${\mathbf{x}}$ at a given time step, a large set of innovations $\left\{\bm{\delta}_{j}\right\}_{j=1\ldots N_{\mathrm{p}}}$ are sampled from the observation error statistics: $\bm{\delta}_{j}\sim N(\mathbf{0},\mathbf{R})$. This yields a set of corresponding incremental updates $\left\{\mathbf{a}_{j}=a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta}_{j})\right\}_{j=1\ldots N_{\mathrm{p}}}$. Since Eq. (7) is only an approximation, $\mathbf{K}({\mathbf{x}})$ is estimated with the least squares problem
$\mathcal{L}_{\mathbf{x}}(\mathbf{K})=\sum_{j=1}^{N_{\mathrm{p}}}\left\|\mathbf{a}_{j}-\bar{\mathbf{a}}-\mathbf{K}({\mathbf{x}})\cdot\left(\bm{\delta}_{j}-\bar{\bm{\delta}}\right)\right\|^{2},$ (9)
where $\bar{\mathbf{a}}=N_{\mathrm{p}}^{-1}\sum_{j=1}^{N_{\mathrm{p}}}\mathbf{a}_{j}$ and $\bar{\bm{\delta}}=N_{\mathrm{p}}^{-1}\sum_{j=1}^{N_{\mathrm{p}}}\bm{\delta}_{j}$.
Next, assuming $\mathbf{R}$ is known, we would like to estimate the analysis $\mathbf{P}^{\mathrm{a}}$ and background $\mathbf{B}$ error covariance matrices, associated to $a_{\bm{\theta}}$ in the Kalman gain interpretation. They depend on ${\mathbf{x}}_{k}$ and hence on $t_{k}$. Given that $\mathbf{K}\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$, it should not be possible to fully determine either $\mathbf{P}^{\mathrm{a}}$ or $\mathbf{B}$, unless $\mathbf{H}$ is invertible. But we can access error covariances in observation space rather than in state space. Within the best linear unbiased estimator framework, it can be shown that
$\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\mathbf{K}\mathbf{R},\qquad\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\left(\mathbf{I}_{\mathrm{y}}-\mathbf{H}\mathbf{K}\right)^{-1}\mathbf{H}\mathbf{K}\mathbf{R}.$ (10)
Hence, one can estimate surrogates of $\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ and of $\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ for the learned $a_{\bm{\theta}}$ from the numerically estimated surrogate $\mathbf{K}$, at any time $t_{k}$. When $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, these boil down to
$\mathbf{P}^{\mathrm{a}}=\mathbf{K},\qquad\mathbf{B}=\left(\mathbf{I}_{\mathrm{x}}-\mathbf{K}\right)^{-1}\mathbf{K}.$ (11)
IV.3 Numerical results
At each $t_{k}$, i.e. over many ${\mathbf{x}}_{k}$ on the attractor of the dynamics, it is possible to estimate, the equivalent $\mathbf{K}({\mathbf{x}}_{k})$, $\mathbf{B}({\mathbf{x}}_{k})$, $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})$ as seen from $a_{\bm{\theta}}$. For the sake of simplicity, $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$ are chosen, such that Eq. (11) stands.
The analysis mean bias $b$ and its standard deviation $s$ are first computed over a long L96 $a_{\bm{\theta}}$-based DA run. We obtain $b\simeq 5.10^{-3}$ and $s\simeq 10^{-3}$ which are indeed very small compared to the typical aRMSE of an either DAN or EnKF run, i.e. $0.20$, meaning that the bias of $a_{\bm{\theta}}$ relative to typical updates is roughly $2.5\%$.
The surrogate $\mathbf{P}^{\mathrm{a}}$, denoted $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and estimated from Eq. (11), is compared to that of a concurrently running well-tuned EnKF with $N_{\mathrm{e}}=40$, whose analysis error covariance matrix is $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ is compared to $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ using a normalized Bures-Wasserstein distance: (Bhatia, Jain, and Lim, 2019)
$d_{\mathrm{BW}}(\mathbf{U},\mathbf{V})=\frac{1}{N_{\mathrm{x}}}\left[\mathrm{Tr}\left\{\mathbf{U}+\mathbf{V}-2\left(\mathbf{V}^{\frac{1}{2}}\mathbf{U}\mathbf{V}^{\frac{1}{2}}\right)^{\frac{1}{2}}\right\}\right]^{\frac{1}{2}},$ (12)
where $\mathbf{U}$ and $\mathbf{V}$ are two semi-definite symmetric matrices. This metric is expected to smoothly account for the unmatched principal axes of $\mathbf{U}$ and $\mathbf{V}$, but also their associated variances (eigenspectra). The time-averaged $d_{\mathrm{BW}}$ distance between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ is $0.013$ whereas it is $0.048$ between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $(0.40)^{2}\mathbf{I}_{\mathrm{x}}$, which approximates $\mathbf{P}^{\mathrm{a}}$ of a well-tuned 3D-Var. Furthermore, the averaged-in-time eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are plotted in Fig. 1. They are remarkably close to each other for the first $10$ modes. Beyond these modes the $a_{\bm{\theta}}$ operator is likely to selectively apply some multiplicative inflation, as one would expect from such stable DA runs.
Figure 1: Time-averaged eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$.
We further compute the principal angles (Bocquet and Carrassi, 2017) of the vector subspaces generated by the $N_{\mathrm{u}}=14$ dominant eigenvalues of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and of $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$. Recall that $N_{\mathrm{u}}=14$ is the dimension of the L96 ${\mathcal{U}}$. The principal angles are intrinsic to the relative position of these subspaces; they do not depend on any coordinate system used to parameterize them. This indicates how close the most unstable directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are in state space. The principal angles are reported in Fig. 2. Overall, the simplex formed by the EnKF is on average the most aligned with ${\mathcal{U}}$.(Bocquet and Carrassi, 2017) The subspace spanned by the dominant axes of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ are also well aligned with ${\mathcal{U}}$, yet progressively diverges when incorporating less unstable directions. For comparison, the principal angles of ${\mathcal{U}}$ with an isotropically randomly sampled $N_{\mathrm{u}}=14$-dimensional subspace are also computed.
Figure 2: Time-averaged principal angles (in degrees) formed by the subspaces spanned by the $N_{\mathrm{u}}=14$ dominant directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ versus ${\mathcal{U}}$, $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ versus ${\mathcal{U}}$, $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ versus $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$, and ${\mathcal{U}}$ versus a randomly sampled $N_{\mathrm{u}}=14$–dimensional subspace.
IV.4 Main interpretation
These numerical results indicate that $a_{\bm{\theta}}$ defined through Eq. (2a), depends on the innovation, but also on the single forecast state when $N_{\mathrm{e}}=1$. This does not hold for the EnKF incremental update which only indirectly depends on the forecast state via the ensemble-based forecast error covariances. Hence, without the need of an ensemble, $a_{\bm{\theta}}$ extracts from the forecast state critical pieces of information on the unstable directions, as shown by the principal angles experiment.
Furthermore, $a_{\bm{\theta}}$ manages to accurately assess the uncertainty attached to these unstable directions as demonstrated by the spectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$. Overall, $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ with $N_{\mathrm{e}}=1$ is on average very close to $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ with $N_{\mathrm{e}}=40$, for the dominant axes, and it applies some inflation onto the less unstable modes as seen by comparing their spectra.(Bocquet, Raanes, and Hannart, 2015) We conclude that $a_{\bm{\theta}}$ directly learns about the dynamics features, as opposed to the regression-based, purely statistical, update in the EnKF.
Essentially, for $a_{\bm{\theta}}$, critical pieces of information of the forecast error covariances of the DA run are encoded, and thus exploitable, in the forecast state alone. From the multiplicative ergodic theorem,(Oseledec, 1968) we know that, in autonomous ergodic dynamical systems such as $\mathcal{M}$, there exists a mapping between each of the system’s states and the corresponding Lyapunov covariant vectors. Further, if the DA run (the forecast and analysis cycle) is considered as an ergodic dynamical system of its own,Carrassi et al. (2008b) the same theorem guarantees the existence of a mapping between the forecast state and the analysis error covariance matrix that $a_{\bm{\theta}}$ guesses. The DA process is not autonomous because it indirectly depends on the truth trajectory, the observation noise, and the observation operators; but a generalized variant of the multiplicative ergodic theorem for non-autonomous random dynamics should be applicable. (Arnold, 1998; Chekroun, Simonnet, and Ghil, 2011; Flandoli and Tonello, 2021; Ghil and Sciamarella, 2023) Hence, we conjecture that $a_{\bm{\theta}}$ must learn such mapping, together with how to process this information and combine it with the innovation.
IV.5 Locality and scalability
Next, we have trained $a_{\bm{\theta}}$ on the L96 model, using the reference configuration, but with a changing state space dimension $N_{\mathrm{x}}$ in between $20$ and $160$. The aRMSEs of well-tuned EnKFs for the changing $N_{\mathrm{x}}$, and picking $N_{\mathrm{e}}=N_{\mathrm{x}}$, has been computed for comparison. The test DAN aRMSEs shows no significant dependence on $N_{\mathrm{x}}$ and are all within $5\%$ of the EnKFs. Hence, because the performance of $a_{\bm{\theta}}$ with an unchanged architecture and number of parameters is barely affected by increasing $N_{\mathrm{x}}$, we conjecture that the learned analysis extracts local pieces of information from the forecast state.
If true, the $a_{\bm{\theta}}$ operator learned for DA on an $N_{\mathrm{x}}=40$ L96 model could be applied directly to an L96 DA run with a different $N_{\mathrm{x}}$. Recall that the L96 states typically exhibit local highs and lows of Rossby-like waves, whose number scales linearly with $N_{\mathrm{x}}$. Thus as long as the spatial extent of those waves is captured by the receptive field of the CNN, the same layers of $a_{\bm{\theta}}$ with the same weights and biases might be able to handle L96 states of distinct dimensionality.
To test this hypothesis, we use the same $a_{\bm{\theta}}$ operator (same weights and biases) learned as before with $N_{\mathrm{e}}=1$, $N_{\mathrm{x}}=40$ but apply it to L96 models with $N_{\mathrm{x}}$ ranging from $20$ to $160$. The corresponding aRMSEs are reported in Fig. 3. This transdimensional transfer works surprisingly well: the aRMSEs are roughly the same for all $N_{\mathrm{x}}$ (between 0.188 and 0.197).
Figure 3: Test aRMSEs of $a_{\bm{\theta}}$ operators learned from either L96 models with varying $N_{\mathrm{x}}$, or learned from the $N_{\mathrm{x}}=40$ L96 model but applied to varying $N_{\mathrm{x}}$ L96 models. The dimension $N_{\mathrm{u}}$ of ${\mathcal{U}}$ is much steeper compared to the slowly increasing aRMSE curves.
This strongly supports the fact that $a_{\bm{\theta}}$ extracts local information from the forecast state (of various dimension in this experiment), relying on its convolution layers. It is therefore able to capture where, in phase-space, the error mass is concentrated. These localized error structures should be related to the localization of the dominant covariant Lyapunov vectors.(Pazó, Rodríguez, and López, 2010; Vannitsem and Lucarini, 2016; Giggins and Gottwald, 2019)
V Conclusions
We have demonstrated on the L96 chaotic model that a learned DL-based analysis $a_{\bm{\theta}}$, key part of a sequential DA scheme, can be almost as accurate as the best possibly tuned EnKF, the benchmark for ensemble filtering methods in this model. More importantly, this learned DA scheme does not require any ensemble and can equally well rely on a single state forecast. Therefore, $a_{\bm{\theta}}$ can presumably read local patterns, that are representative of unstable and uncertain modes, from the forecast state alone. We believe that this is fundamentally made possible by some multiplicative ergodic theorem applied to sequential DA seen as a non-autonomous random dynamical system driven by time-dependent true dynamics and observation operators, and white-in-time observation errors.
To make sure our conclusions were not entirely bound to the L96 model, we carried out a large number of similar experiments on the well-known chaotic Kuramoto-Sivashinski model.(Kuramoto and Tsuzuki, 1976; Sivashinsky, 1977) They all confirm and support these conclusions.
What is achieved by $a_{\bm{\theta}}$ resonates with the parametric EnKF, (Pannekoucke et al., 2016; Pannekoucke, Bocquet, and Ménard, 2018) which encodes the errors of the day in a couple of dynamical ancillary fields, preventing the use of an ensemble. Amazingly, the learned $a_{\bm{\theta}}$ exhibited here is even more radical and extracts that information from the state itself.
Taking a step back, we learned from DL that an accurate and efficient DA analysis operator could capture the dynamical error without an ensemble, leveraging model-specific information. This argues in favor of rethinking popular sequential DA schemes for chaotic dynamics.
Acknowledgements.
The authors acknowledge the support of the project SASIP (grant no. $353$) funded by Schmidt Sciences – a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies. CEREA is a member of Institut Pierre-Simon Laplace (IPSL).
Data Availability Statement
No datasets were used in this article.
Appendix A Architecture of the analysis operator
The learned operator $a_{\bm{\theta}}$ predicts the analysis increment to be added to the forecast/background ensemble, following Eq. (2) The architecture of the network is a basic residual CNN. It begins with an initial convolution that takes $N_{\mathrm{e}}+1$ channels as inputs and, with $N_{\mathrm{f}}$ filters, outputs $N_{\mathrm{f}}$ channels. This initial layer is followed by $N_{\mathrm{b}}$ residual blocks. Each one of these blocks is a succession of $N_{\mathrm{sb}}$ sub-blocks. Each subblock is made of: (i) a convolutional layer with $N_{\mathrm{f}}$ channels as inputs, has $N_{\mathrm{f}}$ filters and a kernel size for each of its filter of $f_{\mathrm{v}}$, (ii) a batch normalization layer, and (iii) an activation function chosen to be mish. (Misra, 2019) The CNN finishes with a final convolutional layer that takes $N_{\mathrm{f}}$ channels as inputs and, with $N_{\mathrm{e}}$ filters, outputs $N_{\mathrm{e}}$ channels. The kernel size of the initial and final channels is $f_{\mathrm{v}}$. A schematic of the CNN architecture is displayed in Fig. 4. The reference configuration, $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$, has about $2\times 10^{5}$ trainable parameters.
Figure 4: Architecture of the residual convolutional network, where $N_{\mathrm{b}}=2$, $N_{\mathrm{sb}}=3$. $\mathrm{conv}_{N_{1},N_{2},f}$ is a generic one-dimensional convolutional layer of dimension $N_{1}$, with $N_{2}$ filters of kernel size $f$.

### v2

@@
-Deep learning-based sequential data assimilation for chaotic dynamics identifies local instabilities from single state forecasts
+Accurate deep learning-based filtering for chaotic dynamics by identifying instabilities without an ensemble
 Abstract
-We investigate the ability to discover data assimilation (DA) schemes meant for chaotic dynamics with deep learning (DL).
+We investigate the ability to discover data assimilation (DA) schemes meant for chaotic dynamics with deep learning.
 The focus is on learning the analysis step of sequential DA, from state trajectories and their observations, using a simple residual convolutional neural network, while assuming the dynamics to be known.
 Experiments are performed with the Lorenz 96 dynamics, which display spatiotemporal chaos and for which solid benchmarks for DA performance exist.
-The accuracy of the states obtained from the learned analysis approaches that of the best possibly tuned ensemble Kalman filter (EnKF), and is far better than that of variational DA alternatives.
+The accuracy of the states obtained from the learned analysis approaches that of the best possibly tuned ensemble Kalman filter, and is far better than that of variational DA alternatives.
 Critically, this can be achieved while propagating even just a single state in the forecast step.
@@
 This is explained by the CNN’s ability to identify local space patterns from this one state, which are used to assess the errors in the analysis.
-This suggests building a new class of efficient DL-based ensemble-free DA algorithms.
+This suggests building a new class of efficient deep learning-based ensemble-free DA algorithms.
 I Introduction
-I.1 Sequential data assimilation and accuracy
-Sequential data assimilation (DA) is a set of algorithms required to accurately track the trajectory of chaotic dynamics, using a dynamical model and imperfect observations.
-It is widely used in numerical weather prediction (NWP), and in many areas of climate sciences, (Carrassi et al., 2018) as a suite of both research and operational tools.
+I.1 Context and problem
+In a simplified but quintessential framework, the goal of data assimilation (DA), and in particular filtering algorithms, is to accurately estimate states ${\mathbf{x}}_{k}^{\mathrm{t}}\in\mathbb{R}^{N_{\mathrm{x}}}$, where “t” stands for truth, at equally spaced times $\tau_{k}$ for $k=0,\ldots,K$ along a trajectory of a dynamical system.
+Hence, they are related by
+${\mathbf{x}}_{k+1}^{\mathrm{t}}={\mathcal{M}}\left({\mathbf{x}}_{k}^{\mathrm{t}}\right),$ (1a)
+where $\mathcal{M}$ is the resolvent over $\tau_{k+1}-\tau_{k}$ of known autonomous, i.e.
+time-independent, dynamics.
+Such goal is achieved from the knowledge of the dynamics $\mathcal{M}$ and of observation vectors $\mathbf{y}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}}$ obtained from the non-accessible states ${\mathbf{x}}_{k}^{\mathrm{t}}$ via observation operators $\mathcal{H}_{k}$, and perturbed by a white-in-time Gaussian noise ${\bm{\varepsilon}}_{k}$ of mean $\mathbf{0}$ and covariance matrix $\mathbf{R}_{k}$:
+$\mathbf{y}_{k}=\mathcal{H}_{k}({\mathbf{x}}_{k}^{\mathrm{t}})+{\bm{\varepsilon}}_{k},\qquad{\bm{\varepsilon}}_{k}\sim N(\mathbf{0},\mathbf{R}_{k}).$ (1b)
+Applied to chaotic hence dynamically unstable dynamics, sequential (in time) algorithms must be used.(Asch, Bocquet, and Nodet, 2016; Carrassi et al., 2018) They alternate an analysis step which, from the newly acquired observation vector $\mathbf{y}_{k}$ in Eq.
+(1b) and the current estimate of the state ${\mathbf{x}}_{k}^{\mathrm{f}}$, provides an updated optimal estimate of the state ${\mathbf{x}}_{k}^{\mathrm{a}}$ called the analysis.
+The subsequent state estimate ${\mathbf{x}}_{k+1}^{\mathrm{f}}$ stems from the forecast step which relies on Eq.
+(1a).
+The estimates in both steps can either be deterministic or probabilistic, often leveraging an ensemble in the latter case.
+Such sequential DA is widely used in numerical weather prediction (NWP), and in many areas of climate sciences,(Carrassi et al., 2018) as a suite of both research and operational tools.
 Classical DA methods are classified into (i) variational methods, such as 3D-Var and 4D-Var, (ii) ensemble-based statistical methods, such as the ensemble Kalman filter (EnKF), and (iii) ensemble variational methods which inherit the assets of the two previous categories.(Asch, Bocquet, and Nodet, 2016) On the one hand, variational methods account for the nonlinearity of models (dynamical model and observation operators), leveraging nonlinear optimization techniques.
@@
 For low-order, chaotic dynamics such as the celebrated Lorenz 96 (L96) model,(Lorenz and Emanuel, 1998) the EnKF significantly outperforms 3D-Var, or a moderately-long window 4D-Var in terms of accuracy, owing to its dynamical representation of the errors.
-This has been emphasized and illustrated in twin experiments.
-(Bocquet and Sakov, 2013) In fact, current implementations of 4D-Var in NWP centers, incorporate a forecast ensemble so as to capture the errors-of-the-day.(Raynaud, Berre, and Desroziers, 2012; Bonavita, Isaksen, and Hólm, 2012) However, in high-dimensional models, these ensemble-based error statistics must necessarily be regularized using techniques known as localization and possibly inflation.(Evensen, 2009)
-I.2 Time-dependent error statistics and dynamical systems
-With a focus on the time-dependent error statistics of sequential DA, it has been conjectured (Carrassi et al., 2008a; Palatella, Carrassi, and Trevisan, 2013) then proven (Gurumoorthy et al., 2017; Bocquet et al., 2017; Crisan and Ghil, 2023) that for linear dynamics and when localization is unnecessary, the forecast and analysis error covariance matrices of the EnKF are confined to the unstable-neutral subspace, denoted as ${\mathcal{U}}$ from now on, of the dynamics.
+This has been emphasized and illustrated in twin experiments.(Bocquet and Sakov, 2013) In fact, current implementations of 4D-Var in NWP centers incorporate a forecast ensemble so as to capture the errors-of-the-day.(Raynaud, Berre, and Desroziers, 2012; Bonavita, Isaksen, and Hólm, 2012) However, in high-dimensional models, these ensemble-based error statistics must necessarily be regularized using techniques known as localization and possibly inflation.(Evensen, 2009) With a focus on the time-dependent error statistics of sequential DA, it has been conjectured (Carrassi et al., 2008a; Palatella, Carrassi, and Trevisan, 2013) then proven (Gurumoorthy et al., 2017; Bocquet et al., 2017; Crisan and Ghil, 2023) that for linear dynamics and when localization is unnecessary, the forecast and analysis error covariance matrices of the EnKF are confined to the unstable-neutral subspace, denoted as ${\mathcal{U}}$ from now on, of the dynamics.
 This subspace is spanned by the covariant Lyapunov vectors associated to non-negative Lyapunov exponents.(Legras and Vautard, 1996) It is precisely when the ensemble size is smaller than the dimension of this subspace that localization is required to avoid divergence of the EnKF.
 Deviating from linear dynamics turns those exact results into approximations, for which these findings were nonetheless numerically confirmed.(Bocquet and Carrassi, 2017; Chen, Carrassi, and Lucarini, 2021; Carrassi et al., 2022)
-I.3 End-to-end data assimilation and deep learning
 This paper focuses on methodological DA and on what can be discovered from deep learning (DL) techniques to improve state-of-the-art DA schemes such as those mentioned above.
@@
 As far as ensemble and Kalman-related DA methods are concerned, it has been proposed to learn their Kalman gain, (Hoang, De Mey, and Talagrand, 1994; Hoang et al., 1998) or parameters thereof, possibly relying on an auto-differentiable implementation of the (En)KF.
-(Haarnoja et al., 2016; Chen, Sanz-Alonso, and Willett, 2022; Luk et al., 2024) A step further, it was also proposed to learn the fully analysis operator using (self-)supervision.(McCabe and Brown, 2021; Boudier et al., 2023) Finally, bypassing the need for dynamical models and DA schemes altogether, DL-based end-to-end methods aim at estimating states of the system from the observations only,(McNally et al., 2024; Vaughan et al., 2024) but with a focus on the feasibility of such endeavor for now.
-I.4 Objective
-In this paper, the forecast model is assumed to be known so as to disentangle intricate interactions of learning simultaneously the DA operators and the dynamics altogether.
-Our objective is to learn the analysis step of a sequential DA scheme meant for chaotic dynamics from a long trajectory of the dynamical system and the associated set of noisy, potentially partial observations.
-Supervised learning is used although self-supervised learning(Bocquet et al., 2019; Brajard et al., 2020; McCabe and Brown, 2021) could have been used instead but is more challenging and rather disconnected from the main results of this paper.
-The (incremental) analysis CNN operator will be referred to as $a_{\bm{\theta}}$, while the full resulting DA scheme will be called DAN.
-We will show that the learned $a_{\bm{\theta}}$ directly discovers and utilizes a fine knowledge of the dynamics, as opposed to agnostic classical DA.
-The nature of these fine dynamical structures learned through $a_{\bm{\theta}}$ will then be discussed and interpreted, together with key implications for the development of modern DA methods.
+(Haarnoja et al., 2016; Chen, Sanz-Alonso, and Willett, 2022; Luk et al., 2024) As a step further, it was also proposed to learn the full analysis operator using (self-)supervision.(McCabe and Brown, 2021; Boudier et al., 2023) Finally, bypassing the need for dynamical models and DA schemes altogether, DL-based end-to-end methods aim at estimating states of the system from the observations only,(McNally et al., 2024; Vaughan et al., 2024) yet so far with a focus on the feasibility of such endeavor.
+I.2 Objectives
+In this paper, the forecast model in Eq. (1a) is assumed to be known so as to avoid intricate interactions when learning the DA operators and the dynamics simultaneously.
+Our objective is to learn the analysis operator of a sequential DA scheme meant for chaotic dynamics from a long trajectory of the dynamical system and the associated set of noisy, possibly sparse observations.
+Hence, it stands out from past studies that exploited DL to learn the dynamics, possibly using DA.
+(Bocquet et al., 2019; Brajard et al., 2020; Bocquet et al., 2020; Brajard et al., 2021; Liu et al., 2022; Wang et al., 2024) The resulting DL-based analysis operator will be referred to as $a_{\bm{\theta}}$, while the full resulting DA scheme will be called DAN.
+We will first explain how to learn such analysis operator from DL.
+It will then be shown numerically that $a_{\bm{\theta}}$ can surprisingly perform as accurately as a well optimized EnKF, even without using an ensemble which strongly contrasts with the common beliefs in methodological DA.
+To interpret this result, we will show using innovative concepts based on a Taylor expansion of the learned $a_{\bm{\theta}}$, that $a_{\bm{\theta}}$ directly discovers and utilizes a fine knowledge of the dynamics, as opposed to agnostic classical DA.
+The nature of these dynamical structures learned through $a_{\bm{\theta}}$ will then be discussed and interpreted.
 II Experimental setup
-We use a twin experiment setup where the states of a trajectory of the dynamics, ${\mathbf{x}}_{k}^{\mathrm{t}}\in\mathbb{R}^{N_{\mathrm{x}}}$ at equally spaced times $t_{k}$ for $k=1\ldots N_{\mathrm{c}}$, are generated from the known dynamics; the superscript “t” stands for truth.
-A set of observation vectors $\mathbf{y}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}}$ are obtained from these states ${\mathbf{x}}_{k}^{\mathrm{t}}$ via an observation operator $\mathcal{H}_{k}$, and perturbed by a white-in-time Gaussian noise ${\bm{\varepsilon}}_{k}$ of mean $\mathbf{0}$ and covariance matrix $\mathbf{R}_{k}\in\mathbb{R}^{N_{\mathrm{y},k}\times N_{\mathrm{y},k}}$:
-$\mathbf{y}_{k}=\mathcal{H}_{k}({\mathbf{x}}_{k}^{\mathrm{t}})+{\bm{\varepsilon}}_{k},\qquad{\bm{\varepsilon}}_{k}\sim N(\mathbf{0},\mathbf{R}_{k}).$ (1)
+With the goal to learn an analysis operator $a_{\bm{\theta}}$ as a key step of a filtering DA scheme for chaotic dynamics, we build a twin experiment within the framework offered by Eqs. (1).
 II.1 Analysis operator and its neural network representation
-Let us define a sequential DA scheme, based on an analysis and forecast (also called background) ensemble.
-The $i$-th members of the analysis and forecast ensembles at time $t_{k}$ are noted ${\mathbf{x}}_{k}^{\mathrm{a},i}$ and ${\mathbf{x}}_{k}^{\mathrm{f},i}$, respectively.
-Denoting ${\mathcal{S}}_{\mathrm{e}}=1\ldots N_{\mathrm{e}}$, the corresponding analysis and forecast ensembles are ${\mathbf{E}}_{k}^{\mathrm{a}}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, ${\mathbf{E}}_{k}^{\mathrm{f}}=\left\{{\mathbf{x}}_{k}^{\mathrm{f},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, respectively, where ${\mathcal{E}}^{\mathrm{e}}=\mathbb{R}^{N_{\mathrm{e}}\times N_{\mathrm{x}}}$.
+Let us define a filtering DA scheme, based on an analysis and forecast ensemble.
+The $i$-th members of the analysis and forecast ensembles at time $\tau_{k}$ are noted ${\mathbf{x}}_{k}^{\mathrm{a},i}$ and ${\mathbf{x}}_{k}^{\mathrm{f},i}$, respectively.
+Denoting ${\mathcal{S}}_{\mathrm{e}}=1,\ldots,N_{\mathrm{e}}$, the corresponding analysis and forecast ensembles are ${\mathbf{E}}_{k}^{\mathrm{a}}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, ${\mathbf{E}}_{k}^{\mathrm{f}}=\left\{{\mathbf{x}}_{k}^{\mathrm{f},i}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}\subset{\mathcal{E}}^{\mathrm{e}}$, respectively, where ${\mathcal{E}}^{\mathrm{e}}=\mathbb{R}^{N_{\mathrm{e}}\times N_{\mathrm{x}}}$.
 The initial ensemble ${\mathbf{E}}_{0}^{\mathrm{f}}$ is obtained from perturbing a random state on the attractor of the dynamics.
-The analysis step of the DA scheme is given by the (incremental) analysis operator $a_{\bm{\theta}}$, which depends on a set of CNN weights and biases, a vector $\bm{\theta}$,
+The analysis step of the DA scheme is given by the (incremental) analysis operator $a_{\bm{\theta}}$, which depends on a set of neural network weights and biases, a vector $\bm{\theta}$,
 ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+a_{\bm{\theta}}\left({\mathbf{E}}_{k}^{\mathrm{f}},\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right),$ (2a)
-where ${\bm{\delta}}_{k}$, the innovation at time $t_{k}$, is defined by
+where ${\bm{\delta}}_{k}$, the innovation at time $\tau_{k}$, is defined by
 ${\bm{\delta}}_{k}\overset{\Delta}{=}\mathbf{y}_{k}-\mathcal{H}_{k}\left({\bar{\mathbf{x}}}_{k}^{\mathrm{f}}\right),\quad{\bar{\mathbf{x}}}_{k}^{\mathrm{f}}\overset{\Delta}{=}\frac{1}{N_{\mathrm{e}}}\sum_{i\in{\mathcal{S}}_{\mathrm{e}}}{\mathbf{x}}_{k}^{\mathrm{f},i}.$ (2b)
 $\mathbf{H}_{k}$ is the tangent linear operator of $\mathcal{H}_{k}$ but any arbitrary injective operator from $\mathbb{R}^{N_{\mathrm{y},k}}$ to $\mathbb{R}^{N_{\mathrm{x}}}$ could be chosen instead.
-The DA forecast step propagates the analysis ensemble,
-${\mathbf{E}}_{k+1}^{\mathrm{f}}={\mathcal{M}}\left({\mathbf{E}}_{k}^{\mathrm{a}}\right),$ (3)
-member-wise; $\mathcal{M}$ is the resolvent over $t_{t+1}-t_{k}$ of the autonomous, i.e. time-independent, dynamics.
-In this paper, we choose $a_{\bm{\theta}}$ to be a simple residual CNN, whose architecture is detailed in Appendix A.
-The internal convolutional layers are chosen to have $N_{\mathrm{f}}$ channels.
-Hence, the internal state of the CNN consists of $N_{\mathrm{f}}$ copies of the latent space which we simply chose to be isomorphic to the state space $\mathbb{R}^{N_{\mathrm{x}}}$.
-Furthermore, the encoder and decoder from state space to latent space and back, are chosen to coincide with the identity.
+The DA forecast step propagates the analysis ensemble, member-wise:
+${\mathbf{E}}_{k+1}^{\mathrm{f}}={\mathcal{M}}\left({\mathbf{E}}_{k}^{\mathrm{a}}\right).$ (3)
+We choose $a_{\bm{\theta}}$ to have a simple residual convolutional neural network (CNN) architecture. A schematic of the CNN architecture is displayed in Fig. 1.
+Figure 1: Architecture of the residual convolutional network, where $N_{\mathrm{b}}=2$, $N_{\mathrm{sb}}=3$.
+$\mathrm{conv}_{N_{1},N_{2},f}$ is a generic one-dimensional convolutional layer of dimension $N_{1}$, with $N_{2}$ filters of kernel size $f$.
+See text for more details.
+It begins with an initial convolution that takes $N_{\mathrm{e}}+1$ channels as inputs and, with $N_{\mathrm{f}}$ filters, outputs $N_{\mathrm{f}}$ channels.
+This initial layer is followed by $N_{\mathrm{b}}$ residual blocks.
+Each one of these blocks is a succession of $N_{\mathrm{sb}}$ sub-blocks.
+Each subblock is made of: (i) a convolutional layer with $N_{\mathrm{f}}$ channels as inputs, which has $N_{\mathrm{f}}$ filters and a kernel size $f_{\mathrm{v}}$ for each of its filter, (ii) a batch normalization layer, and (iii) an activation function chosen to be mish.
+(Misra, 2019) The CNN ends with a final convolutional layer that takes $N_{\mathrm{f}}$ channels as inputs and, with $N_{\mathrm{e}}$ filters, outputs $N_{\mathrm{e}}$ channels.
+The kernel size of the initial and final channels is $f_{\mathrm{v}}$.
+Hence, the internal state of the CNN consists of $N_{\mathrm{f}}$ copies of the latent space which we simply choose to be isomorphic to the state space $\mathbb{R}^{N_{\mathrm{x}}}$.
+Furthermore, the encoder and decoder from state space to latent space and back are chosen to coincide with the identity.
+We have also tested a depth-separable architecture for this residual CNN, with a number of parameters roughly divided by $3$, yielding an accuracy almost as good but longer training times.
+Note that the fundamental results reported in this paper are agnostic to the details of the architecture: this CNN is a mere functional tool to learn an optimal $a_{\bm{\theta}}$.
 II.2 Training scheme
 Towards efficiently learning an optimal $a_{\bm{\theta}}$, we consider $N_{\mathrm{r}}$ such DA runs, based on as many independent concurrent trajectories of the dynamics and as many sequences of observation vectors.
-The DA runs are hence specified by ${\mathbf{E}}_{k}^{\mathrm{a},r}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i,r}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}$ and ${\mathbf{E}}_{k}^{\mathrm{f},r}$ for $r=1\ldots N_{\mathrm{r}}$ and, being iterates through $N_{\mathrm{c}}$ cycles, they depend on $\bm{\theta}$, except for the set of initial conditions ${\mathbf{E}}_{0}^{\mathrm{f},r}$.
+Hence, the DA runs are specified by ${\mathbf{E}}_{k}^{\mathrm{a},r}=\left\{{\mathbf{x}}_{k}^{\mathrm{a},i,r}\right\}_{i\in{\mathcal{S}}_{\mathrm{e}}}$ and ${\mathbf{E}}_{k}^{\mathrm{f},r}$ for $r=1,\ldots,N_{\mathrm{r}}$ and, being iterates through $N_{\mathrm{c}}$ cycles, they depend on $\bm{\theta}$ except for the set of initial conditions ${\mathbf{E}}_{0}^{\mathrm{f},r}$.
 In order to learn an optimal $a_{\bm{\theta}}$, a loss function is defined:
@@
 […]
-It must be pointed out that a proper sequential DA process when seen as a dynamical system, is stable.
+It must be pointed out that a successful sequential DA process, when seen as a dynamical system, is stable.
 (Carrassi et al., 2008b, c) Hence, after a rough starting phase in the training, the learned $a_{\bm{\theta}}$ should yield a numerically stable prediction-assimilation dynamical system.
 In particular, this is likely to avoid exploding gradients.
-By contrast, the task of learning a DL emulator of the dynamics over many consecutive time steps fails because of the unstable nature of the chaotic dynamics.
+By contrast, the task of learning a DL emulator of the dynamics over many consecutive time steps often fails because of the unstable nature of the chaotic dynamics.
 The $N_{\mathrm{r}}$ trajectories are dispatched into a training and validation dataset with a $90\%-10\%$ ratio.
-Moreover, the testing dataset is generated from an independently generated trajectory, long enough to yield converged statistics.
+Overfitting is prevented by an early stopping of the minimization based on the validation score, tantamount to regularization.(Goodfellow, Bengio, and Courville, 2016) Moreover, the testing dataset stems from an independently generated trajectory, long enough to yield converged statistics.
 In the test stage, the DAN scheme is used within a twin experiment using the trajectories and resulting observations of the testing dataset.
-Its performance is assessed using the time-averaged root mean square error (aRMSE) against the truth.
+Its performance is assessed from a single scalar score using the time-averaged root mean square error (aRMSE) of the analysis against the truth:
+$\mathrm{aRMSE}=\frac{1}{K\sqrt{N_{\mathrm{x}}}}\sum_{k=1}^{K}\left\|{\mathbf{x}}_{k}^{\mathrm{t}}-\bar{\mathbf{x}}_{k}^{\mathrm{a}}\right\|,$ (5)
+which, in a cycled DA context, is a reliable indicator of the overall performance of the scheme, whatever its purpose.
 III Numerical results
-The $a_{\bm{\theta}}$ operator is trained on the L96 model, and the results will be interpreted and discussed in the context of this model.
-The L96 model is a one-dimensional model defined over a periodic band of latitude of the Earth atmosphere.
+The $a_{\bm{\theta}}$ operator is trained on the L96 model,(Lorenz and Emanuel, 1998) and the results will be interpreted and discussed in the context of this model.
+L96 is a one-dimensional model defined over a periodic band of latitude of the Earth atmosphere.
 Its ordinary differential equations read
 $\frac{{\mathrm{d}}x_{n}}{{\mathrm{d}}t}=(x_{n+1}-x_{n-2})x_{n-1}-x_{n}+F,$ (6)
-where $x_{N_{x}}=x_{0}$, $x_{-1}=x_{N_{x}-1}$, $x_{-2}=x_{N_{x}-2}$, $F=8$, and $N_{\mathrm{x}}=40$ in the basic configuration.
-The Lyapunov time of this model is $0.60$; it has $13$ positive Lyapunov exponents and a zero Lyapunov exponent as a continuous-in-time autonomous system.
-Hence, the dimension of ${\mathcal{U}}$ is $N_{\mathrm{u}}=14$.
+with $x_{N_{x}}=x_{0}$, $x_{-1}=x_{N_{x}-1}$, $x_{-2}=x_{N_{x}-2}$, $F=8$, and $N_{\mathrm{x}}=40$ in the basic configuration.
+The model has a Lyapunov time of $0.60$.
+It has $13$ positive exponents and, being continuous-in-time and autonomous, it has one zero Lyapunov exponent.
+Hence, the dimension of its unstable-neutral subspace ${\mathcal{U}}$, is $N_{\mathrm{u}}=14$.
 III.1 Hyperparameters sensitivity analysis
 We first carry out a large set of trainings to assess the sensitivity of $a_{\bm{\theta}}$’s performance to its hyperparameters.
-We chose $N_{\mathrm{iter}}=16$, without any significant gain beyond this value while the numerical cost increases due to a deeper backpropagation.
-We assume the model to be fully observed with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, the identity matrix in $\mathbb{R}^{N_{\mathrm{x}}}$, and the observations to be affected by a white-in-time unbiased Gaussian noise of covariance matrix $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, for all time steps.
-This configuration is very often used to benchmark new DA schemes.
+We choose $N_{\mathrm{iter}}=16$, without any significant gain beyond this value while the numerical cost increases due to a deeper backpropagation.
+We first assume the model to be fully observed with $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$, the identity matrix in $\mathbb{R}^{N_{\mathrm{x}}}$, and the observations to be affected by a white-in-time unbiased Gaussian noise of covariance matrix $\mathbf{R}_{k}=\mathbf{I}_{\mathrm{x}}$, for all time steps.
+This configuration is the most widely used to benchmark new DA schemes with L96.
 The ensemble size $N_{\mathrm{e}}$ and the number of filter $N_{\mathrm{f}}$ were selected in a set ranging between $1$ and $40$.
 The number $N_{\mathrm{b}}$ of residual blocks in the CNN and number of subblocks $N_{\mathrm{sb}}$ in each residual block were both chosen in the set $\llbracket 1,6\rrbracket$.
-Because the L96 model has short-range correlations in space, we choose a kernel size of $f_{\mathrm{v}}=5$, even though the CNN receptive field is much larger.
-Overfitting is prevented by an early stopping of the minimization based on the validation score.
+Because L96 has short-range correlations in space, we choose a kernel size of $f_{\mathrm{v}}=5$, even though the CNN receptive field is much larger.
 III.2 First results and robustness
-The test DA runs with the trained $a_{\bm{\theta}}$ are actually all stable in time, yielding an aRMSE significantly below $1$, as expected if DA has any skill over the mere observations.
+The training dataset size per epoch scales linearly with $N_{\mathrm{r}}$, which is chosen to be $2^{18}$ and further discussed in the supplementary material.
+The subsequent test DA runs with the trained $a_{\bm{\theta}}$ are actually all stable in time, yielding an aRMSE significantly below $1$, as expected if DA has any skill over the mere observations.
 Unsurprisingly, we found that the larger the hyperparameters $N_{\mathrm{f}},N_{\mathrm{b}},N_{\mathrm{sb}}$, the smaller the test aRMSEs of the resulting DANs, but that $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$ offer a good compromise for accuracy versus training cost and CNN size.
-This will be the reference configuration.
+This will be the reference configuration, which has about $2\times 10^{5}$ trainable parameters.
 One obvious essential drawback of learning $a_{\bm{\theta}}$ is its non-universality.
@@
 This is a critical research path for end-to-end DA.
-Although not the aim of the present paper, we nonetheless checked the performance of the trained $a_{\bm{\theta}}$, with $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\sigma_{\mathrm{y}}^{2}\mathbf{I}_{\mathrm{x}}$ with $\sigma_{\mathrm{y}}=1$, in test DA runs with similar observations but generated with $\sigma_{\mathrm{y}}$ taking value in between $0.1$ and $3.2$.
+Although not the aim of the present paper, we nonetheless checked the performance of the trained $a_{\bm{\theta}}$, with $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}_{k}=\sigma_{\mathrm{y}}^{2}\mathbf{I}_{\mathrm{x}}$ with $\sigma_{\mathrm{y}}=1$, in test DA runs with similar observations but generated with $\sigma_{\mathrm{y}}$ taking value in between $0.1$ and $3$.
 Yet, in all test runs, DAN remains robust with slightly degraded aRMSEs for $\sigma_{\mathrm{y}}<1$ but aRMSEs at least as good for $\sigma_{\mathrm{y}}>1$, compared to a well-tuned EnKF.
 Well-tuned EnKF always refers here to an EnKF with an ensemble large enough such that localization is unnecessary and relying on the EnKF-N (Bocquet, Raanes, and Hannart, 2015; Raanes, Bocquet, and Carrassi, 2019) to optimally counteract residual sampling errors such that inflation is unnecessary.
-We also learned a unique $a_{\bm{\theta}}$ from observation networks whose density $N_{\mathrm{y}}/N_{\mathrm{x}}$ is randomly and uniformly chosen in the interval $[0,1]$, and $\sigma_{\mathrm{y}}=1$.
-This DAN was then tested on several DA runs, each one with a fixed observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$ taking value in the interval $[0.2,1]$.
+Testing non-trivial $\mathcal{H}_{k}$, we also learned a single $a_{\bm{\theta}}$ from observation networks whose density $N_{\mathrm{y}}/N_{\mathrm{x}}$ is randomly and uniformly chosen in the interval $[0,1]$ at each $\tau_{k}$, and $\sigma_{\mathrm{y}}=1$.
+This DAN was then tested on several DA runs, each one with a constant in time observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$ taking value in the interval $[0.2,1]$.
 In this configuration, $a_{\bm{\theta}}$ performs almost as well or better than well-tuned EnKFs for $0.650.35<N_{\mathrm{y}}/N_{\mathrm{x}}<$ 0.65 and is suboptimal (compared to the EnKF) but still stable outside of this range.
-These results already pleasantly suggest that the findings that follow may remain valid well beyond the specifications of observation operators from which $a_{\bm{\theta}}$ was learned.
+These results already pleasantly suggest that these DL-based DA schemes may remain valid well beyond the specifications of observation operators from which $a_{\bm{\theta}}$ was learned.
+Plots of these experiments and further discussion are provided in the supplementary material.
 III.3 One state forecast
@@
 I, one of the key reason for the success of the EnKF.
-For quantitative comparison, we checked that 3D-Var scores an aRMSE of $0.40$, the best linear filter, i.e.
-$a_{\bm{\theta}}$ learned without activation functions scores $0.38$, whereas well-tuned EnKFs with $N_{\mathrm{e}}=20$ and $N_{\mathrm{e}}=40$ score $0.191$ and $0.179$, respectively.
+For quantitative comparison, we checked that 3D-Var scores an aRMSE of $0.40$, that the best linear filter, i.e.
+$a_{\bm{\theta}}$ learned without activation functions scores $0.384$, whereas well-tuned EnKFs with $N_{\mathrm{e}}=20$ and $N_{\mathrm{e}}=40$ score $0.191$ and $0.179$, respectively.
 Note that the reference $a_{\bm{\theta}}$ but with $N_{\mathrm{f}}=100$ yields an aRMSE of $0.185$, closer to the best EnKF with $N_{\mathrm{e}}=N_{\mathrm{x}}=40$, showing that further improvements are possible even though not the focus of this paper.
-However, the key remark is that a single state forecast, $N_{\mathrm{e}}=1$ in $a_{\bm{\theta}}$, is as efficient as using a large ensemble.
+These key aRMSE scores are arranged in a table in the supplementary material.
+However, the pivotal remark is that a single state forecast, $N_{\mathrm{e}}=1$ in $a_{\bm{\theta}}$, is as efficient as using a large ensemble.
 Furthermore, the need for localization and inflation is completely obviated.
@@
 all channels’ last layer feature maps converge to the same state.
-It is likely that a better local minimum of the loss could be obtained with complex encoder and decoder (Peyron et al., 2021) and infusing diversity in the NN through Monte Carlo dropouts,(McCabe and Brown, 2021) so as to obtain an $a_{\bm{\theta}}$ leveraging the ensemble.
+It is likely that a better local minimum of the loss could be obtained with complex encoder and decoder (Peyron et al., 2021) and infusing diversity in the CNN through Monte Carlo dropouts,(McCabe and Brown, 2021) so as to obtain an $a_{\bm{\theta}}$ leveraging the ensemble.
 Nonetheless, the local minimum reached in our trainings, yield an accuracy with $N_{\mathrm{e}}=1$ worthy of a well-tuned EnKF.
-That is why we shall concentrate in the following on interpreting this astonishing result, especially in Sec.
-IV, using dynamical systems theory.
-Consequently, the analysis operator is hereafter learned in the reference configuration but with $N_{\mathrm{e}}=1$.
-IV Interpretation and supporting results
-In this section, we focus on the main result: a learned DA method with a single state forecast achieves performance on par with a well tuned EnKF.
-We would like to understand the reason for this performance and what $a_{\bm{\theta}}$ is learning.
-IV.1 Kalman gain-based expansion
-Towards this goal, we build an incremental classical Kalman update (Kalman, 1960; Ghil and Malanotte-Rizzoli, 1991) that would be a good match to $a_{\bm{\theta}}$ seen as a mathematical map, at least for small analysis increments.
+That is why we shall concentrate in the following on interpreting this astonishing result for which we shall use, especially in Sec.
+IV, dynamical systems theory.
+Therefore, the analysis operator is hereafter learned in the reference configuration but with $N_{\mathrm{e}}=1$.
+IV Interpretation
+In this section, we focus on the remarkable finding that a learned DA method with a single state $N_{\mathrm{e}}=1$ forecast achieves performance on par with a well-tuned EnKF.
+We wish to understand the reason for this performance by investigating what $a_{\bm{\theta}}$ learns.
+To that end, an innovative expansion of $a_{\bm{\theta}}$ in terms of more familiar DA operators is carried out.
+IV.1 Operator expansion of $a_{\bm{\theta}}$
+Towards this goal, we look for a classical Kalman update (Kalman, 1960; Ghil and Malanotte-Rizzoli, 1991) that would be a good match to $a_{\bm{\theta}}$ seen as a mathematical map, at least for small analysis increments.
 The first diagnostic is the mean anomaly generated by $a_{\bm{\theta}}$, i.e.
@@
 This is an Ansatz of $a_{\bm{\theta}}$ where $\mathbf{K}({\mathbf{x}})\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$ is meant to stand as a Kalman gain surrogate.
-With the propagation of a single state, classical sequential DA methods would typically resemble 3D-Var, and the gain would not depend on the forecast state (the first input variable of $a_{\bm{\theta}}$).
-The 3D-Var equivalent gain would be
-$\mathbf{K}_{\mathrm{3D-Var}}=\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\left(\mathbf{R}+\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\right)^{-1},$ (8)
-where $\mathbf{R}$ and/or $\mathbf{H}$ might depend on time if the monitoring system does so, and where $\mathbf{B}$ represents some expected climatology of the DA errors.
-Interestingly, we also learned $a_{\bm{\theta}}$ replacing Eq.
-(2a) with the simplified ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+a_{\bm{\theta}}\left(\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right)$, whereby losing $a_{\bm{\theta}}$’s ability to extract information from ${\mathbf{E}}_{k}^{\mathrm{f}}$.
-This yields an aRMSE of $0.38$ in test DA runs, unsurprisingly close to the $0.40$ of our 3D-Var.
-Hence, learning an optimal constant-in-time $\mathbf{K}$ of an (En)KF(Luk et al., 2024), a configuration subsumed by this specific $a_{\bm{\theta}}$, is significantly suboptimal in this context.
-IV.2 Estimation of the Kalman gain and error covariances
-Once $a_{\bm{\theta}}$ has been obtained from training, and considering a fixed background state ${\mathbf{x}}$ at a given time step, a large set of innovations $\left\{\bm{\delta}_{j}\right\}_{j=1\ldots N_{\mathrm{p}}}$ are sampled from the observation error statistics: $\bm{\delta}_{j}\sim N(\mathbf{0},\mathbf{R})$.
-This yields a set of corresponding incremental updates $\left\{\mathbf{a}_{j}=a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta}_{j})\right\}_{j=1\ldots N_{\mathrm{p}}}$.
+By contrast, with the propagation of a single state, classical sequential DA methods would typically resemble 3D-Var, and the gain would not depend on the forecast state (the first input variable of $a_{\bm{\theta}}$).
+Interestingly, we also learned a simplified ${\hat{a}}_{\bm{\theta}}$ replacing Eq.
+(2a) with ${\mathbf{E}}_{k}^{\mathrm{a}}={\mathbf{E}}_{k}^{\mathrm{f}}+{\hat{a}}_{\bm{\theta}}\left(\mathbf{H}_{k}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}_{k}{\bm{\delta}}_{k}\right)$, whereby losing $a_{\bm{\theta}}$’s ability to extract information from ${\mathbf{E}}_{k}^{\mathrm{f}}$, similarly to 3D-Var.
+This yields an aRMSE of $0.382$ in test DA runs, unsurprisingly close to the $0.40$ of our 3D-Var.
+Hence, learning an optimal constant-in-time $\mathbf{K}$ of an (En)KF,(Luk et al., 2024) a configuration subsumed by this specific ${\hat{a}}_{\bm{\theta}}$, is significantly suboptimal in this context.
+IV.2 Identifying the operators in this expansion
+Once $a_{\bm{\theta}}$ has been obtained from training, and considering a fixed forecast state ${\mathbf{x}}$ at a given time step, a large set of innovations $\left\{\bm{\delta}_{j}\right\}_{j=1,\ldots,N_{\mathrm{p}}}$ are sampled from the observation error statistics: $\bm{\delta}_{j}\sim N(\mathbf{0},\mathbf{R})$.
+This yields a set of corresponding incremental updates $\left\{\mathbf{a}_{j}=a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta}_{j})\right\}_{j=1,\ldots,N_{\mathrm{p}}}$.
 Since Eq.
@@
 where $\bar{\mathbf{a}}=N_{\mathrm{p}}^{-1}\sum_{j=1}^{N_{\mathrm{p}}}\mathbf{a}_{j}$ and $\bar{\bm{\delta}}=N_{\mathrm{p}}^{-1}\sum_{j=1}^{N_{\mathrm{p}}}\bm{\delta}_{j}$.
-Next, assuming $\mathbf{R}$ is known, we would like to estimate the analysis $\mathbf{P}^{\mathrm{a}}$ and background $\mathbf{B}$ error covariance matrices, associated to $a_{\bm{\theta}}$ in the Kalman gain interpretation.
-They depend on ${\mathbf{x}}_{k}$ and hence on $t_{k}$.
-Given that $\mathbf{K}\in\mathbb{R}^{N_{\mathrm{x}}\times N_{\mathrm{y}}}$, it should not be possible to fully determine either $\mathbf{P}^{\mathrm{a}}$ or $\mathbf{B}$, unless $\mathbf{H}$ is invertible.
-But we can access error covariances in observation space rather than in state space.
-Within the best linear unbiased estimator framework, it can be shown that
-$\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\mathbf{K}\mathbf{R},\qquad\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}=\left(\mathbf{I}_{\mathrm{y}}-\mathbf{H}\mathbf{K}\right)^{-1}\mathbf{H}\mathbf{K}\mathbf{R}.$ (10)
-Hence, one can estimate surrogates of $\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ and of $\mathbf{H}\mathbf{B}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}$ for the learned $a_{\bm{\theta}}$ from the numerically estimated surrogate $\mathbf{K}$, at any time $t_{k}$.
-When $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$ and $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$, these boil down to
-$\mathbf{P}^{\mathrm{a}}=\mathbf{K},\qquad\mathbf{B}=\left(\mathbf{I}_{\mathrm{x}}-\mathbf{K}\right)^{-1}\mathbf{K}.$ (11)
-IV.3 Numerical results
-At each $t_{k}$, i.e.
-over many ${\mathbf{x}}_{k}$ on the attractor of the dynamics, it is possible to estimate, the equivalent $\mathbf{K}({\mathbf{x}}_{k})$, $\mathbf{B}({\mathbf{x}}_{k})$, $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})$ as seen from $a_{\bm{\theta}}$.
-For the sake of simplicity, $\mathcal{H}=\mathbf{I}_{\mathrm{x}}$, $\mathbf{R}=\mathbf{I}_{\mathrm{x}}$ are chosen, such that Eq.
-(11) stands.
+Next, assuming $\mathbf{R}$ is known, we would like to estimate the analysis error covariance matrix $\mathbf{P}^{\mathrm{a}}$ associated to $a_{\bm{\theta}}$ in the Kalman gain expansion.
+It depends on ${\mathbf{x}}_{k}$ and hence on $\tau_{k}$.
+Within the best linear unbiased estimator framework, $\mathbf{K}$ is related to $\mathbf{P}^{\mathrm{a}}$ through $\mathbf{K}=\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}$, so that from Eq.
+(8):
+$a_{\bm{\theta}}({\mathbf{x}},\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta})\approx\mathbf{P}^{\mathrm{a}}\mathbf{H}^{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}\mathbf{R}^{-1}\bm{\delta},$ (10)
+which suggests that an expansion in the second variable ${\bm{\zeta}}\in\mathbb{R}^{N_{\mathrm{x}}}$ of $a_{\bm{\theta}}$ yields
+$a_{\bm{\theta}}({\mathbf{x}},{\bm{\zeta}})\approx\mathbf{P}^{\mathrm{a}}({\mathbf{x}})\cdot{\bm{\zeta}}.$ (11)
+Hence, $\mathbf{P}^{\mathrm{a}}$ can be estimated using Eq.
+(11) either from a least squares loss similar to Eq.
+(9) or from the Jacobian of $a_{\bm{\theta}}$ with respect to ${\bm{\zeta}}$ leveraging auto-differentiable DL libraries.
+IV.3 What is learned? - Supporting numerical results
+At each $\tau_{k}$, i.e.
+over many ${\mathbf{x}}_{k}$ on the forecast model’s attractor, it is possible to estimate $\mathbf{K}({\mathbf{x}}_{k})$ and $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})$ from the expansion of $a_{\bm{\theta}}$.
+For the sake of simplicity, $\mathcal{H}_{k}=\mathbf{I}_{\mathrm{x}}$, $\mathbf{R}_{k}=\mathbf{I}_{\mathrm{x}}$, in which case $\mathbf{P}^{\mathrm{a}}({\mathbf{x}}_{k})=\mathbf{K}({\mathbf{x}}_{k})$.
 The analysis mean bias $b$ and its standard deviation $s$ are first computed over a long L96 $a_{\bm{\theta}}$-based DA run.
@@
 The surrogate $\mathbf{P}^{\mathrm{a}}$, denoted $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and estimated from Eq.
-(11), is compared to that of a concurrently running well-tuned EnKF with $N_{\mathrm{e}}=40$, whose analysis error covariance matrix is $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$.
+(11), is compared to that of a concurrent well-tuned EnKF with $N_{\mathrm{e}}=40$, whose analysis error covariance matrix is $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$.
 $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ is compared to $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ using a normalized Bures-Wasserstein distance: (Bhatia, Jain, and Lim, 2019)
@@
 The time-averaged $d_{\mathrm{BW}}$ distance between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ is $0.013$ whereas it is $0.048$ between $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $(0.40)^{2}\mathbf{I}_{\mathrm{x}}$, which approximates $\mathbf{P}^{\mathrm{a}}$ of a well-tuned 3D-Var.
-Furthermore, the averaged-in-time eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are plotted in Fig.
-1.
+The time-averaged eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are plotted in Fig.
+2.
 They are remarkably close to each other for the first $10$ modes.
-Beyond these modes the $a_{\bm{\theta}}$ operator is likely to selectively apply some multiplicative inflation, as one would expect from such stable DA runs.
+Beyond these modes the $a_{\bm{\theta}}$ operator is likely to selectively apply some (multiplicative) inflation, as one would expect from such stable DA runs.
 Figure 2: Time-averaged eigenspectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$.
 We further compute the principal angles(Bocquet and Carrassi, 2017) of the vector subspaces generated by the $N_{\mathrm{u}}=14$ dominant eigenvalues of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and of $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$.
+They are reported in Fig.
+3.
 Recall that $N_{\mathrm{u}}=14$ is the dimension of the L96 ${\mathcal{U}}$.
@@
 This indicates how close the most unstable directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ and $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ are in state space.
-The principal angles are reported in Fig.
-2.
-Overall, the simplex formed by the EnKF is on average the most aligned with ${\mathcal{U}}$.(Bocquet and Carrassi, 2017) The subspace spanned by the dominant axes of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ are also well aligned with ${\mathcal{U}}$, yet progressively diverges when incorporating less unstable directions.
-For comparison, the principal angles of ${\mathcal{U}}$ with an isotropically randomly sampled $N_{\mathrm{u}}=14$-dimensional subspace are also computed.
+From Fig.
+3, we observe that the simplex formed by the EnKF is on average the most aligned with ${\mathcal{U}}$.(Bocquet and Carrassi, 2017) The subspace spanned by the dominant axes of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ are also well aligned with ${\mathcal{U}}$, yet progressively diverges when incorporating less unstable directions.
+For comparison, the principal angles of ${\mathcal{U}}$ with an isotropically randomly sampled $N_{\mathrm{u}}=14$-dimensional subspace are also plotted in Fig.
+3.
 Figure 3: Time-averaged principal angles (in degrees) formed by the subspaces spanned by the $N_{\mathrm{u}}=14$ dominant directions of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$ versus ${\mathcal{U}}$, $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ versus ${\mathcal{U}}$, $\mathbf{P}^{\mathrm{a}}_{\mathrm{EnKF}}$ versus $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$, and ${\mathcal{U}}$ versus a randomly sampled $N_{\mathrm{u}}=14$–dimensional subspace.
@@
 This does not hold for the EnKF incremental update which only indirectly depends on the forecast state via the ensemble-based forecast error covariances.
-Hence, without the need of an ensemble, $a_{\bm{\theta}}$ extracts from the forecast state critical pieces of information on the unstable directions, as shown by the principal angles experiment.
+Hence, without the need for an ensemble, $a_{\bm{\theta}}$ extracts from the forecast state critical pieces of information on the unstable directions, as shown by the principal angles experiment.
 Furthermore, $a_{\bm{\theta}}$ manages to accurately assess the uncertainty attached to these unstable directions as demonstrated by the spectra of $\mathbf{P}^{\mathrm{a}}_{\mathrm{DAN}}$.
@@
 IV.5 Locality and scalability
-Next, we have trained $a_{\bm{\theta}}$ on the L96 model, using the reference configuration, but with a changing state space dimension $N_{\mathrm{x}}$ in between $20$ and $160$.
+Next, we have trained $a_{\bm{\theta}}$ on the L96 model using the reference configuration with $N_{\mathrm{e}}=1$, but with a changing state space dimension $N_{\mathrm{x}}$ in between $20$ and $160$.
 The aRMSEs of well-tuned EnKFs for the changing $N_{\mathrm{x}}$, and picking $N_{\mathrm{e}}=N_{\mathrm{x}}$, has been computed for comparison.
 The test DAN aRMSEs shows no significant dependence on $N_{\mathrm{x}}$ and are all within $5\%$ of the EnKFs.
-Hence, because the performance of $a_{\bm{\theta}}$ with an unchanged architecture and number of parameters is barely affected by increasing $N_{\mathrm{x}}$, we conjecture that the learned analysis extracts local pieces of information from the forecast state.
+Hence, because the performance of $a_{\bm{\theta}}$ with an unchanged architecture and the same number of parameters is barely affected by increasing $N_{\mathrm{x}}$, we conjecture that the learned analysis extracts local pieces of information from the forecast state.
 If true, the $a_{\bm{\theta}}$ operator learned for DA on an $N_{\mathrm{x}}=40$ L96 model could be applied directly to an L96 DA run with a different $N_{\mathrm{x}}$.
-Recall that the L96 states typically exhibit local highs and lows of Rossby-like waves, whose number scales linearly with $N_{\mathrm{x}}$.
+Recall that the L96 states exhibit local highs and lows of Rossby-like waves, whose number scales linearly with $N_{\mathrm{x}}$.
 Thus as long as the spatial extent of those waves is captured by the receptive field of the CNN, the same layers of $a_{\bm{\theta}}$ with the same weights and biases might be able to handle L96 states of distinct dimensionality.
-To test this hypothesis, we use the same $a_{\bm{\theta}}$ operator (same weights and biases) learned as before with $N_{\mathrm{e}}=1$, $N_{\mathrm{x}}=40$ but apply it to L96 models with $N_{\mathrm{x}}$ ranging from $20$ to $160$.
+To test this hypothesis, we use the same $a_{\bm{\theta}}$ operator (same weights and biases) learned as before with $N_{\mathrm{e}}=1$, $N_{\mathrm{x}}=40$ but apply it now to L96 models with $N_{\mathrm{x}}$ ranging from $20$ to $160$.
 The corresponding aRMSEs are reported in Fig.
-3.
-This transdimensional transfer works surprisingly well: the aRMSEs are roughly the same for all $N_{\mathrm{x}}$ (between 0.188 and 0.197).
-Figure 3: Test aRMSEs of $a_{\bm{\theta}}$ operators learned from either L96 models with varying $N_{\mathrm{x}}$, or learned from the $N_{\mathrm{x}}=40$ L96 model but applied to varying $N_{\mathrm{x}}$ L96 models.
-The dimension $N_{\mathrm{u}}$ of ${\mathcal{U}}$ is much steeper compared to the slowly increasing aRMSE curves.
-This strongly supports the fact that $a_{\bm{\theta}}$ extracts local information from the forecast state (of various dimension in this experiment), relying on its convolution layers.
+4, which shows that these aRMSEs are roughly the same for all $N_{\mathrm{x}}$ (between 0.188 and 0.197).
+This demonstrates that this transdimensional transfer works surprisingly well.
+Figure 4: Test aRMSEs (blue and yellow full lines) of $a_{\bm{\theta}}$ operators learned from either L96 models with varying $N_{\mathrm{x}}$, or learned from the $N_{\mathrm{x}}=40$ L96 model but applied to varying $N_{\mathrm{x}}$ L96 models.
+The dimension $N_{\mathrm{u}}$ of ${\mathcal{U}}$ (gray dashed line) is much steeper compared to the slowly increasing aRMSE curves.
+This strongly supports the fact that $a_{\bm{\theta}}$ extracts local information from the forecast state (of various dimensions in this experiment), relying on its convolution layers.
 It is therefore able to capture where, in phase-space, the error mass is concentrated.
-These localized error structures should be related to the localization of the dominant covariant Lyapunov vectors.(Pazó, Rodríguez, and López, 2010; Vannitsem and Lucarini, 2016; Giggins and Gottwald, 2019)
+We hypothesize that these localized error structures are related to the localization of the dominant covariant Lyapunov vectors.(Pazó, Rodríguez, and López, 2010; Vannitsem and Lucarini, 2016; Giggins and Gottwald, 2019) A proper mathematical definition of such spatial localization can be found in these references.
 V Conclusions
-We have demonstrated on the L96 chaotic model that a learned DL-based analysis $a_{\bm{\theta}}$, key part of a sequential DA scheme, can be almost as accurate as the best possibly tuned EnKF, the benchmark for ensemble filtering methods in this model.
+Using the L96 chaotic model we have demonstrated that a learned DL-based analysis $a_{\bm{\theta}}$, key part of a sequential DA (often referred to as a filtering scheme) can be almost as accurate as the best possibly tuned EnKF, the benchmark for ensemble filtering methods in this model.
 More importantly, this learned DA scheme does not require any ensemble and can equally well rely on a single state forecast.
-Therefore, $a_{\bm{\theta}}$ can presumably read local patterns, that are representative of unstable and uncertain modes, from the forecast state alone.
+Therefore, $a_{\bm{\theta}}$ appears to be able to retrieve local patterns, representative of unstable and uncertain modes, from the forecast state alone.
 We believe that this is fundamentally made possible by some multiplicative ergodic theorem applied to sequential DA seen as a non-autonomous random dynamical system driven by time-dependent true dynamics and observation operators, and white-in-time observation errors.
@@
 What is achieved by $a_{\bm{\theta}}$ resonates with the parametric EnKF, (Pannekoucke et al., 2016; Pannekoucke, Bocquet, and Ménard, 2018) which encodes the errors of the day in a couple of dynamical ancillary fields, preventing the use of an ensemble.
-Amazingly, the learned $a_{\bm{\theta}}$ exhibited here is even more radical and extracts that information from the state itself.
+Amazingly, our learned $a_{\bm{\theta}}$ is even more radical and extracts that information from the state itself.
 Taking a step back, we learned from DL that an accurate and efficient DA analysis operator could capture the dynamical error without an ensemble, leveraging model-specific information.
-This argues in favor of rethinking popular sequential DA schemes for chaotic dynamics.
+This promotes a rethinking of the popular sequential DA schemes for chaotic dynamics.
+Supplementary material
+See the supplementary material for further details on the evaluation of the $a_{\bm{\theta}}$-based DA method, as reported in Sec.
+III.
+Specifically, a table of the key aRMSE scores mentioned Sec.
+III is provided, as well as plots of the aRMSE curves related to the sensitivity experiments mentioned in Sec.
+III.
 Acknowledgements.
+The authors are grateful to the Editor Jürgen Kurths and an anonymous Reviewer for their comments and suggestions on the original version of the manuscript.
 The authors acknowledge the support of the project SASIP (grant no.
-$353$) funded by Schmidt Sciences – a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies.
+$G-24-66154$) funded by Schmidt Sciences – a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies.
 CEREA is a member of Institut Pierre-Simon Laplace (IPSL).
@@
 No datasets were used in this article.
-Appendix A Architecture of the analysis operator
-The learned operator $a_{\bm{\theta}}$ predicts the analysis increment to be added to the forecast/background ensemble, following Eq.
-(2) The architecture of the network is a basic residual CNN.
-It begins with an initial convolution that takes $N_{\mathrm{e}}+1$ channels as inputs and, with $N_{\mathrm{f}}$ filters, outputs $N_{\mathrm{f}}$ channels.
-This initial layer is followed by $N_{\mathrm{b}}$ residual blocks.
-Each one of these blocks is a succession of $N_{\mathrm{sb}}$ sub-blocks.
-Each subblock is made of: (i) a convolutional layer with $N_{\mathrm{f}}$ channels as inputs, has $N_{\mathrm{f}}$ filters and a kernel size for each of its filter of $f_{\mathrm{v}}$, (ii) a batch normalization layer, and (iii) an activation function chosen to be mish.
-(Misra, 2019) The CNN finishes with a final convolutional layer that takes $N_{\mathrm{f}}$ channels as inputs and, with $N_{\mathrm{e}}$ filters, outputs $N_{\mathrm{e}}$ channels.
-The kernel size of the initial and final channels is $f_{\mathrm{v}}$.
-A schematic of the CNN architecture is displayed in Fig.
-4.
-The reference configuration, $N_{\mathrm{f}}=40$, $N_{\mathrm{b}}=5$, and $N_{\mathrm{sb}}=5$, has about $2\times 10^{5}$ trainable parameters.
-Figure 4: Architecture of the residual convolutional network, where $N_{\mathrm{b}}=2$, $N_{\mathrm{sb}}=3$.
-$\mathrm{conv}_{N_{1},N_{2},f}$ is a generic one-dimensional convolutional layer of dimension $N_{1}$, with $N_{2}$ filters of kernel size $f$.
+Supplementary Material: Accurate deep learning-based filtering for chaotic dynamics by identifying instabilities without an ensemble
+In the following DA, DL, CNN, EnKF, RMSE and aRMSE are the initialisms and acronyms of data assimilation, deep learning, convolutional neural network, ensemble Kalman filter, root mean square error (of the analysis state versus the true state), and time-averaged root mean square error (of the analysis state versus the true state), respectively.
+All the $a_{\bm{\theta}}$ operators trained and tested here are based on $N_{\mathrm{e}}=1$, i.e.
+no ensemble but a single state forecast.
+V.1 Impact of the training dataset size on the DA tests
+A series of $a_{\bm{\theta}}$ operators are learned from increasingly larger training and validation datasets (per epoch), which is achieved by increasing the number of trajectories $N_{\mathrm{r}}$ in the datasets in the range $[2^{10},2^{22}]$, and displayed in Fig.
+5.
+The hyperparameters of the CNN architecture are those of the reference configuration.
+Figure 5: Test aRMSEs of $a_{\bm{\theta}}$ operators learned from a L96 models with a varying number of trajectories $N_{\mathrm{r}}$ in the training dataset, commensurate to the size of the training dataset per epoch.
+The batch size, i.e.
+the number of trajectories in the training batch, is chosen in this experiment to be $\mathrm{min}\left(N_{\mathrm{r}}/10,2^{11}=2,048\right)$ in order to be able to maintain the targeted training and validation datasets ratio, while seeking a large batch for numerical efficiency.
+As a compromise between accuracy and computational time for this study, most of the experiments are conducted with $N_{\mathrm{r}}=2^{18}=262,144$.
+V.2 Impact of the number of filters of $a_{\bm{\theta}}$ as a CNN on the DA tests
+A series of $a_{\bm{\theta}}$ operators are learned from an increasingly larger number of channels/filters $N_{\mathrm{f}}$.
+All other hyperparameters of the CNN architecture are those of the reference configuration.
+Figure 6: Test aRMSEs of the DAN method applied to the L96 model, as a function of the number of channels/filters $N_{\mathrm{f}}$ in $a_{\bm{\theta}}$.
+As a compromise between accuracy and computational time for this paper, most of the experiments are conducted with $N_{\mathrm{f}}=40$.
+V.3 Impact of the observation noise magnitude on the DA tests
+The impact of observation noise in the DA test experiments with magnitude distinct from the noise magnitude used in the training of $a_{\bm{\theta}}$ is discussed in the paper in Section III.B.
+Figure 7 displays the corresponding DAN and EnKF aRMSE curves.
+In all cases the analysis aRMSE remains below the corresponding observational error standard deviation $\sigma_{\mathrm{y}}$.
+Figure 7: Test aRMSEs of the DAN method, and of the EnKF $N_{\mathrm{e}}=20,40$ applied to the L96 model, as a function of the true standard deviation of the assimilated observations.
+V.4 Impact of the sparsity of the observations on the DA tests
+The impact on the test aRMSEs of the DAN scheme with non-trivial observation operators $\mathcal{H}_{k}$ is discussed in the paper in Section III.B, with an experiment where $a_{\bm{\theta}}$ is trained with varying $\mathcal{H}_{k}$ of time-dependent random sparsity ratios, and then tested with observation operators of fixed in time sparsity ratio.
+Figure 8 displays the corresponding DAN and EnKF aRMSE curves.
+The observation error standard deviation is $\sigma_{\mathrm{y}}=1$.
+Figure 8: Test aRMSEs of the DAN method, and of the EnKF $N_{\mathrm{e}}=20,40$ applied to the L96 model, as a function of the observation density $N_{\mathrm{y}}/N_{\mathrm{x}}$.
+Moreover, we report here (though not in the paper) that, as a sanity check, we learned $a_{\bm{\theta}}$ and tested the resulting DAN scheme with a staggered observation operator, where only the variables of the odd sites of the L96 model are observed (a classical test in methodological DA studies).
+We obtain for the DL-based DA scheme and a well-tuned EnKF with $N_{\mathrm{e}}=40$, aRMSEs of $0.289$ and $0.288$, respectively, which suggests that non-trivial observation operators are properly dealt with by $a_{\bm{\theta}}$.
+V.5 Key aRMSE scores of the main DA methods
+For comparison and for reference, the aRMSE scores of the classical DA methods and of the DL-based DAN schemes tested in this paper, are compiled in Tab.
+1.
+The tested well-tuned EnKFs are based on the finite-size EnKF (EnKF-N) DA method with has implicit adaptive inflation and relies here on an ensemble size of either $N_{\mathrm{e}}=20$ or $N_{\mathrm{e}}=40$, beyond the unstable-neutral subspace dimension $N_{\mathrm{u}}=14$ where localization is unnecessary.
+The architecture and hyperparameters of the DL-based DA schemes are those of the reference configuration described in the paper, with possibly the exception of $N_{\mathrm{e}}$ and $N_{\mathrm{f}}$ which are hence explicitly specified in the table.
+Table 1: Compilation of the key aRMSE scores of the classical and DL-based DA methods tested in this paper.
+DA method well-tuned classical DL-based aRMSE
+EnKF-N, $N_{\mathrm{e}}=20$ yes $0.191$
+EnKF-N, $N_{\mathrm{e}}=40$ yes $0.179$
+3D-Var yes $0.40$
+$a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.191$
+$a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=100$ yes $0.185$
+linear $a_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.384$
+simplified ${\hat{a}}_{\bm{\theta}}$, $N_{\mathrm{e}}=1$, $N_{\mathrm{f}}=40$ yes $0.382$
+V.6 Typical RMSE time series of DAN and of a well-tuned EnKF
+Figure 9 illustrates time series of instant RMSEs for the DAN scheme in the reference configuration with $N_{\mathrm{e}}=1$, and for a well-tuned EnKF with $N_{\mathrm{e}}=40$, actually the EnKF-N.
+These time series include the initial spin-up time period.
+Note that the EnKF starts from a perturbed true state, whereas DAN starts from a random state on the L96 attractor.
+Figure 9: Instant RMSEs of the DAN scheme ($N_{\mathrm{e}}=1$) and of a well-tuned EnKF ($N_{\mathrm{e}}=40$) as a function of the L96 model time.